4x + 5 = -40
Subtract 5 from each side.
4x=-45
4(-10)= -40
4(-1)= -4
4(-1/4)= -1
-40+(-4)+(-1)=
-40-4-1= -45
-45+5(from the orginal question)= -40
x= 10 1/4= 10.25

Found 2 solutions by Nate, veghead:
Answer by Nate(3500)

4x = -45
x = -11.25

Answer by veghead(2)

4x + 5 = -40
Subtract 5 from each side for now.
4x=-45
4(-10)= -40
4(-1)= -4
4(-1/4)= -1
-40+(-4)+(-1)=
-40-4-1= -45
-45+5(from the orginal question)= -40
x= 10 1/4= 10.25